Using the letters of the word 'EDUCATION' how many words using 6 letters can be made so
that every word contains at least 4 vowels?

Grade:11

2 Answers

Ritesh Khatri
6 years ago
vowels = {EUAIO}
consonents = {DCTN}
 
words with Atleat 4 vowels = {exactly 4 vowels and 2 consonents} + {exactly 5 vovels and 1 consonent}
5C4*4C2 + 5C5*4C1
=5*6 + 1*4
=30 + 4
=34
 
Waris
5 years ago
vowels = {EUAIO}consonents = {DCTN} words with Atleat 4 vowels = {exactly 4 vowels and 2 consonents} + {exactly 5 vovels and 1 consonent}= 5C5*4C1*6factorial + 5C5*4C*6factorial= 24480In case one combination selects all the combination of 4 vowels and combination of 2 consonents then we use permutation to get all the shuffled 6 letter word with 4 vowels and 2 consonentsSimilarly case 2